# AT 2025abao: the fourth luminous red nova in M 31

### v1

@@
We present photometric and spectroscopic observations of the luminous red nova (LRN) AT 2025abao, the fourth discovered in M 31. The LRN, associated with the AGB star WNTR23bzdiq, was discovered during the fast rise following the minimum phase. It reached the peak at mag ( mag), and then it settled onto a long-duration plateau in the red bands, lasting 70 days, while it was slowly linearly declining in the blue bands. The object showed similarities at peak with the canonical LRNe V838 Monocerotis, V1309 Scorpii, and with the faint and fast-evolving AT 2019zhd, the third LRN in M31, though the later evolution is different. Spectroscopically, AT 2025abao evolved as a canonical LRN: the early spectra present a blue continuum with narrow Balmer lines in emission; at peak, the spectral continuum has cooled to a yellow colour, with a photospheric temperature of 6000 K. Balmer lines have weakened while absorption lines from metals (Fe i, Fe ii, Sc ii, Ba ii, Ti ii) have developed, and in particular broad (FWHM km s-1) from the UV Ca ii H&K lines. Medium- and high-resolution spectra reveal a counter-P Cygni absorption profile in H. Finally, late time spectra show an orange continuum ( K), a return in strength of the Balmer lines and the formation of molecular absorption bands. AT 2025abao is the rare case of a LRN with detailed archival information regarding the progenitor system. For the first time, we obtained the spectral energy distribution in the infrared of the precursor of a LRN, which is consistent with that of an M giant/AGB. We propose that the dichotomy of light curve behaviour in LRNe (two peaks vs. plateau) can be explained by the extent and H-richness of the common envelope.
@@
Luminous Red Novae (LRNe; Kulkarni et al. 2007) are astrophysical transients likely resulting from the coalescence of two non-degenerate stars due to the loss of systemic angular momentum following the ejection of a common envelope (e.g., Tylenda et al., 2011; Ivanova et al., 2013b, a; Kochanek et al., 2014; Pejcha, 2014; MacLeod et al., 2022). They are a subclass of “interacting gap transients” (IGTs; Pastorello and Fraser, 2019; Cai et al., 2022b; Reguitti et al., 2026). IGTs have absolute magnitudes at maximum in the range mag, and fill the observational gap that separates the brightest novae from the faintest supernovae. In addition, their photometric and spectroscopic evolution is dominated by the effects of the interaction with a circumstellar medium (CSM, see Pastorello et al. 2019b for a review on the phenomenon).
LRNe are identified through a few recurrent observational characteristics. Firstly, they present a light curve with a double peak. The first peak has a short duration (1 week); it is typically very luminous and characterised by a blue colour. This first maximum is followed a few weeks (or even months) later by a long-lasting second peak or a plateau-like phase, which is redder and usually fainter than the first111However, see the case of AT 2011kp (a.k.a. NGC4490-OT2011; Pastorello et al., 2019b; Smith et al., 2016), which had a second peak that was more luminous than the first one..
@@
For a number of extragalactic LRNe, the progenitor systems were detected in pre-outburst archival images from deep ground-based facilities or space telescopes (Fraser et al., 2011; Mauerhan et al., 2015; Smith et al., 2016; Blagorodnova et al., 2017; Cai et al., 2019; Pastorello et al., 2021c; Blagorodnova et al., 2021; Cai et al., 2022a). From these findings, it was discovered that the typical progenitors of LRNe are yellow hypergiant stars, with masses up to 50 (Guidolin et al., 2026), and are suggested to sit in the Hertzsprung gap of the HR diagram (Tranin et al., 2025; Addison et al., 2022).
Most LRNe are discovered in distant galaxies, hence they are usually faint targets, making challenging their observational campaigns and, consequently, limiting the amount of information we can infer from their study. For this reason, it is important to find and study IGTs in nearby galaxies, especially within the Local Group. A handful of LRNe were observed in the Milky Way: V4332 Sgr (Martini et al., 1999; Tylenda et al., 2005), OGLE-2002-BLG360 (Tylenda et al., 2013), V838 Mon (Munari et al., 2002; Bond et al., 2003; Tylenda, 2005), and V1309 Sco (Mason et al., 2010; Tylenda et al., 2011). Three others were instead observed in M 31: M31-RV (Mould et al., 1990; Boschi and Munari, 2004), M31-LRN2015, (Kurtenkov et al., 2015; Williams et al., 2015; MacLeod et al., 2017; Lipunov et al., 2017; Blagorodnova et al., 2020), and AT 2019zhd (Pastorello et al., 2021b).
@@ 2 Discovery
AT 2025abao was officially discovered on 2025 October 19.8196333Universal Time is used throughout the paper. (Modified Julian Date, or MJD = 60967.8196) by the Mobile Astronomical System of Telescope-Robots program (Kechin et al., 2025), at an unfiltered magnitude of 17.3 (Vega mag), and immediately reported to the Transient Name Server444https://www.wis-tns.org/ (TNS). However, the first advisory of a brightening from a known transient was made by Koichi Itagaki on October 17.4798, at an unfiltered magnitude of 17.8 (Vegamag)555http://www.cbat.eps.harvard.edu/unconf/followups/J00384865+4046079.html. An even earlier pre-discovery measurement was provided by Fabregat et al. (2025) on October 14. The LRN appeared in the south-west quadrant of the very nearby Andromeda galaxy (M 31), at celestial coordinates 00:38:48.62, 40:46:07.6 ( West and South of the nucleus), in the outskirts of its spiral disk. The location of the transient is shown in Figure 1. AT 2025abao was correctly classified on 2025 November 2 as a LRN by Taguchi and Maeda (2025a), who noted similarities with another LRN found in M 31, AT 2019zhd (Pastorello et al., 2021b). For the distance to M 31, we assume the same value adopted by Pastorello et al. (2021b) ( mag, Mpc). Its redshift is (Falco et al., 1999). However, we note that the peaks of the spectral lines of AT 2025abao are blueshifted by . The additional radial velocity is likely caused by the rotation curve of the galaxy disk, with the object being located on the side approaching us at 150 km s-1 (Chemin et al., 2009). The Galactic reddening towards AT 2025abao is modest, although not negligible, mag (Schlafly and Finkbeiner, 2011). Given the peripheral location of AT 2025abao in the disk of M 31, additional contribution to the dust extinction is likely negligible. Hereafter, in agreement with K25, we will only account for the dust attenuation within the Milky Way.
3 Photometric evolution of AT 2025abao
@@ 3.1 Observations and data reduction
While K25 presented early photometric data of AT 2025abao, this work complements their light curve with previously unpublished pre-LRN photometric points from the Zwicky Transient Factory (ZTF; Bellm et al., 2019; Graham et al., 2019). The initial phases following the LRN outburst were exclusively monitored by amateur astronomers and all-sky surveys, such as ZTF and the Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018). Our multi-band optical and near-infrared (NIR) follow-up campaign began shortly after the spectroscopic classification of the object, and lasted for approximately three months. The Swift space telescope was also triggered to get ultraviolet (UV) photometry at early phases. The facilities used to collect our photometric data are listed in Table 1. The photometric data were reduced using standard procedures with the dedicated ECsnoopy pipeline666ECsnoopy is a package for SN photometry using PSF fitting and/or template subtraction developed by E. Cappellaro. A package description can be found at https://sngroup.oapd.inaf.it/ecsnoopy.html.. Instrumental magnitudes were determined through the Point Spread Function (PSF)-fitting technique, and calibrated against the catalogue of the Sloan Digital Sky Survey (SDSS) survey (Ahumada et al., 2020). The Swift/UVOT UV data were reduced with the HEASOFT pipeline v. 6.35.2777NASA High Energy Astrophysics Science Archive Research Center – Heasarc 2014., by aperture photometry within a fixed radius of 5”. We also retrieved calibrated forced photometry measurements from the ALeRCE Broker (Förster et al., 2021) for ZTF, and from the public Forced Photometry server888https://fallingstar-data.com/forcedphot/ (Shingles et al., 2021) for ATLAS. Given its apparent brightness, the object was also observed extensively by amateur astronomers999More observations of AT 2025abao during the outburst are presented by Mikolajczyk et al. (2026).; the magnitudes were measured via aperture photometry on their pre-reduced frames (corrected for bias, flat-field, and dark frames), and calibrated their measurements against the AAVSO Photometric All-Sky Survey101010https://www.aavso.org/apass catalogue (Henden, 2019). The final UV, optical (Sloan, Johnson, and ATLAS), and NIR magnitudes are listed in the CDS, while the light curves are plotted in Fig. 2.
3.2 Light Curve of AT 2025abao, and its interpretation
@@ 3.2.1 Pre-outburst phase
The pre-outburst evolution of AT 2025abao, from the slow luminosity rise to a broad peak, the superposed magnitude fluctuations, and the following luminosity decline, closely resembles that observed in V1309 Sco (Tylenda et al., 2011). In both cases, the peculiar light-curve evolution is likely a consequence of unstable mass transfer between the stellar components of a binary system (Roche Lobe overflow; see, e.g., Pejcha, 2014; Pejcha et al., 2016b). This culminates with the formation of a common envelope embedding the stellar system which becomes progressively optically thick (Ivanova et al., 2013b). A slow pre-LRN luminosity rise was also observed in other LRNe discovered in M 31, such as M31-LRN-2015 (Blagorodnova et al., 2020) and AT 2019zhd (Pastorello et al., 2021a), as well as in several extra-galactic LRNe, although with some limitations due to a lower monitoring cadence and the poorer signal-to-noise datasets111111The slow brightenings heralding LRN eruptions were observed for AT 2014ib (SNhunt248; Kankare et al., 2015; Mauerhan et al., 2015), AT 2015dl (M101-2015OT1; Blagorodnova et al., 2017; Goranskij et al., 2016), AT 2020hat (Pastorello et al., 2021c; Reguitti et al., 2026), AT 2021blu (Pastorello et al., 2023) and, more marginally, in AT 2011kp (Pastorello et al., 2019a) and AT 2021biy (Cai et al., 2022a)..
@@ 3.2.2 LRN outburst
In the case of AT 2025abao, this phase is followed by a fast brightening of about five magnitudes in the -band (six in ): this is the onset of the LRN outburst. The rapid rise lasts about one week from the discovery of AT 2025abao. Then, its light curve settles on a sort of plateau phase during which the object maintains a nearly constant brightness ( mag). To constrain the timing of the main LRN maximum, we perform a 2nd-order polynomial fit to the -band data collected between MJD = 60973 and MJD = 61000, determining that the light curve peaked on MJD 60986.30.2, at a magnitude of mag. This epoch will be used as a reference for all phases considered in this paper. We note that the bluer-band light curves peak earlier ( d in , d in ), while a photometric peak is not well-defined in the red bands, where a plateau is noticeable, lasting for about two months. This plateau is followed by a sudden decline of the light curve. We measured the post-maximum decline rates in the optical bands by fitting the light curve in each filter with a straight line. The decline is steepest in the -band ( mag 100 d-1), and less and less steep as we move towards redder filters (, , mag 100 d-1). The -band light curve instead is flat, and the -band one is actually still rising, up to phase d, when it reached its maximum.
Individual patterns of LRN light curves are not fully understood. However, there is a general consensus that the characteristic double-peaked shape is due to a merging event. As a consequence of the common envelope ejection and the loss of systemic angular momentum, the two components of the binary system may merge. If this happens, the main LRN outburst is produced by the violent ejection of material following the coalescence of the two stellar cores, and its subsequent interaction with the pre-existing common envelope (see, e.g., Metzger and Pejcha, 2017; MacLeod et al., 2017, 2022; Matsumoto and Metzger, 2022; Kirilov et al., 2025).
@@ 3.4 Comparisons with similar objects
As comparison objects, we select a few faint LRNe well monitored during the slow pre-outburst rise phase. Given that AT 2025abao has an extensive and well-sampled light curve since a few years before the LRN event (K25), our primary reference object is V1309 Sco (Mason et al., 2010; Tylenda et al., 2011), although the two objects have significantly different light curves during the outburst. AT 2025abao reaches an absolute magnitude at the -band peak of mag (Fig. 4), hence it lies at the faint edge of the extragalactic LRNe brightness distribution, although it is significantly more luminous than Galactic objects such as V1309 Sco and V4332 Sgr (Pastorello et al., 2019b). AT 2019zhd (in M 31; Pastorello et al., 2021b) and the Galactic V838 Mon (Goranskij et al., 2020, and references therein) have similar absolute magnitude at maximum as AT 2025abao, and are hence selected as comparison objects.
@@
Due to the plateau-like feature after the first peak, the light curve of AT 2025abao is reminiscent of other LRNe that do not exhibit an evident second red peak, instead showing a long phase of nearly constant luminosity. Known LRNe with plateau-like light curves are AT 2021afy (Pastorello et al., 2023), although it is 4.5 mag more luminous and at the bright edge of the LRNe brightness distribution; AT 2018bwo, whose plateau reaches mag (Blagorodnova et al., 2021; Pastorello et al., 2023); AT 2020hat, that peaks at mag (Pastorello et al., 2021c). However, we remark that AT 2018bwo was discovered already after the first peak due to the Solar conjunction, hence the object could have reached a higher peak luminosity.

### v2

@@
We present photometric and spectroscopic observations of the luminous red nova (LRN) AT 2025abao, the fourth discovered in M 31. The LRN, associated with the AGB star WNTR23bzdiq, was discovered during the fast rise following the minimum phase. It reached the peak at mag ( mag), and then it settled onto a long-duration plateau in the red bands, lasting 70 days, while it was slowly linearly declining in the blue bands. The object showed similarities at peak with the canonical LRNe V838 Monocerotis, V1309 Scorpii, and with the faint and fast-evolving AT 2019zhd, the third LRN in M31, though the later evolution is different. Spectroscopically, AT 2025abao evolved as a canonical LRN: the early spectra present a blue continuum with narrow Balmer lines in emission; at peak, the spectral continuum has cooled to a yellow colour, with a photospheric temperature of 6000 K. Balmer lines have weakened while absorption lines from metals (Fe i, Fe ii, Sc ii, Ba ii, Ti ii) have developed, and in particular broad (FWHM700 km s-1) from the UV Ca ii H&K lines. Medium- and high-resolution spectra reveal a counter-P Cygni absorption profile in H. Finally, late time spectra show an orange continuum (4000-5000 K), a return in strength of the Balmer lines and the formation of molecular absorption bands. AT 2025abao is the rare case of a LRN with detailed archival information regarding the progenitor system. For the first time, we obtained the spectral energy distribution in the infrared of the precursor of a LRN, which is consistent with that of an M giant/AGB. We propose that the dichotomy of light curve behaviour in LRNe (two peaks vs. plateau) can be explained by the extent and H-richness of the common envelope.
@@
Luminous Red Novae (LRNe; Kulkarni et al. 2007) are astrophysical transients likely resulting from the coalescence of two non-degenerate stars due to the loss of systemic angular momentum following the ejection of a common envelope (e.g., Tylenda et al., 2011; Ivanova et al., 2013b, a; Kochanek et al., 2014; Pejcha, 2014; MacLeod et al., 2022). They are a subclass of “interacting gap transients” (IGTs; Pastorello and Fraser, 2019; Cai et al., 2022b; Reguitti et al., 2026). IGTs have absolute magnitudes at maximum in the range mag, and fill the observational gap that separates the brightest novae from the faintest supernovae. In addition, their photometric and spectroscopic evolution is dominated by the effects of the interaction with a circumstellar medium (CSM, see Pastorello et al. 2019 for a review on the phenomenon).
LRNe are identified through a few recurrent observational characteristics. Firstly, they present a light curve with a double peak. The first peak has a short duration (1 week); it is typically very luminous and characterised by a blue colour. This first maximum is followed a few weeks (or even months) later by a long-lasting second peak or a plateau-like phase, which is redder and usually fainter than the first111However, see the case of AT 2011kp (a.k.a. NGC4490-OT2011; Pastorello et al., 2019; Smith et al., 2016), which had a second peak that was more luminous than the first one..
@@
For a number of extragalactic LRNe, the progenitor systems were detected in pre-outburst archival images from deep ground-based facilities or space telescopes (Fraser et al., 2011; Mauerhan et al., 2015; Smith et al., 2016; Blagorodnova et al., 2017; Cai et al., 2019; Pastorello et al., 2021b; Blagorodnova et al., 2021; Cai et al., 2022a). From these findings, it was discovered that the typical progenitors of LRNe are yellow hypergiant stars, with masses up to 50 (Guidolin et al., 2026), and are suggested to sit in the Hertzsprung gap of the HR diagram (Tranin et al., 2025; Addison et al., 2022).
Most LRNe are discovered in distant galaxies, hence they are usually faint targets, making challenging their observational campaigns and, consequently, limiting the amount of information we can infer from their study. For this reason, it is important to find and study IGTs in nearby galaxies, especially within the Local Group. A handful of LRNe were observed in the Milky Way: V4332 Sgr (Martini et al., 1999; Tylenda et al., 2005), OGLE-2002-BLG360 (Tylenda et al., 2013), V838 Mon (Munari et al., 2002; Bond et al., 2003; Tylenda, 2005), and V1309 Sco (Mason et al., 2010; Tylenda et al., 2011). Three others were instead observed in M 31: M31-RV (Mould et al., 1990; Boschi and Munari, 2004), M31-LRN2015, (Kurtenkov et al., 2015; Williams et al., 2015; MacLeod et al., 2017; Lipunov et al., 2017; Blagorodnova et al., 2020), and AT 2019zhd (Pastorello et al., 2021a).
@@ 2 Discovery
AT 2025abao was officially discovered on 2025 October 19.8196333Universal Time is used throughout the paper. (Modified Julian Date, or MJD = 60967.8196) by the Mobile Astronomical System of Telescope-Robots program (Kechin et al., 2025), at an unfiltered magnitude of 17.3 (Vega mag), and immediately reported to the Transient Name Server444https://www.wis-tns.org/ (TNS). However, the first advisory of a brightening from a known transient was made by Koichi Itagaki on October 17.4798, at an unfiltered magnitude of 17.8 (Vegamag)555http://www.cbat.eps.harvard.edu/unconf/followups/J00384865+4046079.html. An even earlier pre-discovery measurement was provided by Fabregat et al. (2025) on October 14. The LRN appeared in the south-west quadrant of the very nearby Andromeda galaxy (M 31), at celestial coordinates 00:38:48.62, 40:46:07.6 ( West and South of the nucleus), in the outskirts of its spiral disk. The location of the transient is shown in Figure 1. AT 2025abao was correctly classified on 2025 November 2 as a LRN by Taguchi and Maeda (2025a), who noted similarities with another LRN found in M 31, AT 2019zhd (Pastorello et al., 2021a). For the distance to M 31, we assume the same value adopted by Pastorello et al. (2021a) ( mag, Mpc). Its redshift is (Falco et al., 1999). However, we note that the peaks of the spectral lines of AT 2025abao are blueshifted by . The additional radial velocity is likely caused by the rotation curve of the galaxy disk, with the object being located on the side approaching us at 150 km s-1 (Chemin et al., 2009). The Galactic reddening towards AT 2025abao is modest, although not negligible, mag (Schlafly and Finkbeiner, 2011). Given the peripheral location of AT 2025abao in the disk of M 31, additional contribution to the dust extinction is likely negligible. Hereafter, in agreement with K25, we will only account for the dust attenuation within the Milky Way.
3 Photometric evolution
@@ 3.1 Observations and data reduction
While K25 presented early photometric data of AT 2025abao, this work complements their light curve with previously unpublished pre-LRN photometric points from the Zwicky Transient Factory (ZTF; Bellm et al., 2019; Graham et al., 2019). The initial phases following the LRN outburst were exclusively monitored by amateur astronomers and all-sky surveys, such as ZTF and the Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018). Our multi-band optical and near-infrared (NIR) follow-up campaign began shortly after the spectroscopic classification of the object, and lasted for approximately four months. The Swift space telescope was also triggered to get ultraviolet (UV) photometry at early phases. The facilities used to collect our photometric data are listed in Table 1. The photometric data were reduced using standard procedures with the dedicated ECsnoopy pipeline666ECsnoopy is a package for SN photometry using PSF fitting and/or template subtraction developed by E. Cappellaro. A package description can be found at https://sngroup.oapd.inaf.it/ecsnoopy.html.. Instrumental magnitudes were determined through the Point Spread Function (PSF)-fitting technique, and calibrated against the catalogue of the Sloan Digital Sky Survey (SDSS) survey (Ahumada et al., 2020). The Swift/UVOT UV data were reduced with the HEASOFT pipeline v. 6.35.2777NASA High Energy Astrophysics Science Archive Research Center – Heasarc 2014., by aperture photometry within a fixed radius of 5”. We also retrieved calibrated forced photometry measurements from the ALeRCE Broker (Förster et al., 2021) for ZTF, and from the public Forced Photometry server888https://fallingstar-data.com/forcedphot/ (Shingles et al., 2021) for ATLAS. Given its apparent brightness, the object was also observed extensively by amateur astronomers999More observations of AT 2025abao during the outburst are presented by Mikolajczyk et al. (2026).; the magnitudes were measured via aperture photometry on their pre-reduced frames (corrected for bias, flat-field, and dark frames), and calibrated their measurements against the AAVSO Photometric All-Sky Survey101010https://www.aavso.org/apass catalogue (Henden, 2019). The final UV, optical (Sloan, Johnson, and ATLAS), and NIR magnitudes are listed in the CDS, while the light curves are plotted in Fig. 2.
3.2 Light curve, and its interpretation
@@ 3.2.1 Pre-outburst phase
The pre-outburst evolution of AT 2025abao, from the slow luminosity rise to a broad peak, the superposed magnitude fluctuations, and the following luminosity decline, closely resembles that observed in V1309 Sco (Tylenda et al., 2011). In both cases, the peculiar light-curve evolution is likely a consequence of unstable mass transfer between the stellar components of a binary system (Roche Lobe overflow; see, e.g., Pejcha, 2014; Pejcha et al., 2016b). This culminates with the formation of a common envelope embedding the stellar system which becomes progressively optically thick (Ivanova et al., 2013b). A slow pre-LRN luminosity rise was also observed in other LRNe discovered in M 31, such as M31-LRN-2015 (Blagorodnova et al., 2020) and AT 2019zhd (Pastorello et al., 2021a), as well as in several extra-galactic LRNe, although with some limitations due to a lower monitoring cadence and the poorer signal-to-noise datasets111111The slow brightenings heralding LRN eruptions were observed for AT 2014ib (SNhunt248; Kankare et al., 2015; Mauerhan et al., 2015), AT 2015dl (M101-2015OT1; Blagorodnova et al., 2017; Goranskij et al., 2016), AT 2020hat (Pastorello et al., 2021b; Reguitti et al., 2026), AT 2021blu (Pastorello et al., 2023) and, more marginally, in AT 2011kp (Pastorello et al., 2019) and AT 2021biy (Cai et al., 2022a)..
@@ 3.2.2 LRN outburst
In the case of AT 2025abao, this phase is followed by a fast brightening of about five magnitudes in the -band (six in ): this is the onset of the LRN outburst. The rapid rise lasts about one week from the discovery of AT 2025abao. Then, its light curve settles on a sort of plateau phase during which the object maintains a nearly constant brightness ( mag). To constrain the timing of the main LRN maximum, we perform a 2nd-order polynomial fit to the -band data collected between MJD = 60973 and MJD = 61000, determining that the light curve peaked on MJD 60986.30.2, at a magnitude of mag. This epoch will be used as a reference for all phases considered in this paper. We note that the bluer-band light curves peak earlier ( d in , d in ), while a photometric peak is not well-defined in the red bands, where a plateau is noticeable, lasting for about two months. We measured the post-maximum decline rates in the optical bands by fitting the light curve in each filter with a straight line. The decline is steepest in the -band ( mag 100 d-1), and less and less steep as we move towards redder filters (, , mag 100 d-1). The -band light curve instead is flat, and the -band one is actually still rising, up to phase d, when it reached its maximum. The plateau phase ends at d and is followed by a sudden fall of the light curve, and then by a linear decline.
Individual patterns of LRN light curves are not fully understood. However, there is a general consensus that the characteristic double-peaked shape is due to a merging event (Banerjee et al., 2026). As a consequence of the common envelope ejection and the loss of systemic angular momentum, the two components of the binary system may merge. If this happens, the main LRN outburst is produced by the violent ejection of material following the coalescence of the two stellar cores, and its subsequent interaction with the pre-existing common envelope (see, e.g., Metzger and Pejcha, 2017; MacLeod et al., 2017, 2022; Matsumoto and Metzger, 2022; Kirilov et al., 2025).
@@ 3.4 Comparisons with similar objects
As comparison objects, we select a few faint LRNe well monitored during the slow pre-outburst rise phase. Given that AT 2025abao has an extensive and well-sampled light curve since a few years before the LRN event (K25), our primary reference object is V1309 Sco (Mason et al., 2010; Tylenda et al., 2011), although the two objects have significantly different light curves during the outburst. AT 2025abao reaches an absolute magnitude at the -band peak of mag (Fig. 4), hence it lies at the faint edge of the extragalactic LRNe brightness distribution, although it is significantly more luminous than Galactic objects such as V1309 Sco and V4332 Sgr (Pastorello et al., 2019). AT 2019zhd (in M 31; Pastorello et al., 2021a) and the Galactic V838 Mon (Goranskij et al., 2020, and references therein) have similar absolute magnitude at maximum as AT 2025abao, and are hence selected as comparison objects.
@@
Due to the plateau-like feature after the first peak, the light curve of AT 2025abao is reminiscent of other LRNe that do not exhibit an evident second red peak, instead showing a long phase of nearly constant luminosity. Known LRNe with plateau-like light curves are AT 2021afy (Pastorello et al., 2023), although it is 4.5 mag more luminous and at the bright edge of the LRNe brightness distribution; AT 2018bwo, whose plateau reaches mag (Blagorodnova et al., 2021; Pastorello et al., 2023); AT 2020hat, that peaks at mag (Pastorello et al., 2021b). However, we remark that AT 2018bwo was discovered already after the first peak due to the Solar conjunction, hence the object could have reached a higher peak luminosity.